OpenStudy (anonymous):

find the slope (−5, −1) and (2, −3)

OpenStudy (lgbasallote):

go istim!

OpenStudy (istim):

slope=y/x. Or, m=y2-y1/x2-x1.

OpenStudy (istim):

IsTim is active.

OpenStudy (anonymous):

??????

OpenStudy (istim):

Ok, you know x,y right.

OpenStudy (anonymous):

yes

OpenStudy (istim):

Then treat the first values at (x1,y1) and the second value as (y2,x2). Then just plug it into the formula I gave.

OpenStudy (anonymous):

ok how would i write that .....it wouldn't be (3,-2.5) would it???

OpenStudy (anonymous):

i am not sure

OpenStudy (istim):

No. It wouldn't. Here. Look at this example.